# Genetic Instruct: Scaling up Synthetic Generation of Coding Instructions for Large Language Models

### v1

@@ Abstract
Large Language Models (LLMs) rely on instruction samples for alignment, but creating these datasets poses challenges, particularly in expert-dependent tasks like coding, which can be cost-prohibitive. One approach to mitigate these challenges is synthesizing data using another LLM. In this paper, we introduce a scalable method for generating synthetic instructions to enhance the code generation capability of LLMs. The proposed algorithm, Genetic-Instruct, mimics evolutionary processes, utilizing self-instruction to create numerous synthetic samples from a limited number of seeds. Genetic-Instruct is designed for efficient scaling of the generation process. Fine-tuning multiple coding LLMs with the synthetic samples demonstrates a significant improvement in their code generation accuracy compared to the baselines.
spacing=nonfrench
Genetic Instruct: Scaling up Synthetic Generation of
Coding Instructions for Large Language Models
Somshubra Majumdar*, Vahid Noroozi thanks: Equal contribution, Sean Narenthiran, Aleksander Ficek, Jagadeesh Balam, Boris Ginsburg NVIDIA Santa Clara, CA 15213, USA {smajumdar,vnoroozi,snarenthiran}@nvidia.com
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated significant advancements in programming and coding problems and can be readily used as code assistant models (Liang et al., 2024). To fully harness their capabilities, a process known as alignment (Ouyang et al., 2022) is necessary, which relies on paired instruction-solution examples to guide the model’s behavior and responses. One challenge in alignment is creating diverse and complex instructions, which can be costly, particularly in domains like coding that require specialized expertise. One strategy to address the scarcity of instructional examples for training is to generate synthetic instructions using another LLM. Research has shown that synthetic instructions can be effective not only for coding applications (Luo et al., 2024; Rozière et al., 2024; Yu et al., 2024) but also for general tasks (Wang et al., 2023; Honovich et al., 2023; Xu et al., 2024). This approach can be seen as a form of knowledge distillation or transfer from a more capable LLM to another.
@@
Figure 1: Genetic-Instruct Process. A small seed population is operated on by an Instructor-LLM to perform crossover/mutation operations. A Coder-LLM then generates an appropriate code solution, which is judged for correctness by a Judge-LLM. Once a required population size is reached, the samples are de-duplicated and form the final population.
In this paper, we introduce a scalable algorithm named Genetic-Instruct, illustrated in Figure 1. Drawing inspiration from evolutionary algorithms, Genetic-Instruct is designed to generate synthetic instructions for coding problems. The algorithm begins with a small set of high-quality initial instruction samples and employs some LLMs to generate instructions and their corresponding code outputs through two evolutionary operations: crossover and mutation. The crossover operation is adapted from self-instruct (Wang et al., 2023), which prompts an LLM to generate new examples based on the context of a few-shot examples. The LLM generates new samples using seed instructions as a foundation, akin to a crossover operation. In the mutation operation, an LLM is given an instruction and asked to evolve it into another instruction based on predefined rules (Luo et al., 2024). Subsequently, another LLM proficient in coding generates the output or solution to the instruction, typically comprising coding solutions.
We also introduce a fitness function that employs an LLM to assess the correctness and quality of the generated instruction and its solution. Samples that pass these evaluations and checks are added to the population pool, and the evolutionary process continues until the target population size is reached. The entire pipeline is designed for efficient parallel execution with multiple colonies of populations. We extend previous work by providing a solution that enables existing operators with scalability as needed.
We generate a large dataset of synthetic coding instructions using the proposed Genetic-Instruct algorithm, starting from a limited set of coding problems. We then assessed the effectiveness of this dataset by training multiple open-source software (OSS) LLMs. Our work is dedicated to enhancing OSS models, and we strictly avoid using any closed-source data or models. The coding models trained with our synthetic dataset were evaluated across a range of coding benchmarks, showcasing the superior performance of our approach compared to other algorithms designed for generating synthetic coding instructions.
@@ 2 Previous Works
Synthetic data generation has emerged as a viable alternative to the time-intensive and expensive process of collecting data from human interactions with LLM systems. One notable method for automated data generation is Self-Instruct (Wang et al., 2023), which employs a pre-trained base LLM to generate instruction-output pairs from a small seed dataset. The resulting dataset is then used to fine-tune the base LLM, resulting in an Instruction-tuned LLM. This work is not focused on coding problems and generate instruction for general tasks.
A limitation of Self-Instruct is that the synthesized samples are often simple and lack diversity, failing to generate challenging instructions that require additional steps to arrive at the solution. To address this, Xu et al. (2024) introduced Evol-Instruct, an enhancement of Self-Instruct that integrates the concept of mutation of instructions to create more complex tasks. By introducing meta-instructions that increase the reasoning depth, add constraints to solutions, or evolve instructions with similar concepts (in-breadth evolution), Evol-Instruct produces a diverse set of complex instructions that significantly improve the performance of the instruction-tuned model. This concept was further adapted for the coding domain by Luo et al. (2024) in a variant called WizardCoder, where models trained on these datasets exhibit enhanced coding capabilities. Our work extends WizardCoder mutation by supporting mutating multiple generations of the same sample and contrarily open-sourcing our work.
While Self-Instruct and Evol-Instruct primarily focus on synthesizing code generation-style instruction-solution pairs, Yu et al. (2024) introduced the CodeOcean dataset, which adopts a novel approach to synthesizing complex tasks for code models. The models trained on CodeOcean, referred to as WaveCoder, are adept at code completion, documentation, fixing, and translation. Contrary to previous methods that construct instructions based on in-context seed instructions, the CodeOcean dataset leverages code snippets from existing codebases as context. The LLM is then tasked with constructing a valid instruction, one of the four tasks, followed by an appropriate solution, leading to a more diverse and challenging set of instructions. Other works such as PromptBreeder (Fernando et al., 2023) combine mutation and crossover but at the prompt level and then operating on the provided instruction to create a new solution. Our work mutates instructions by performing Wizard-Coder (and potentially Wave-Coder) on the given instruction as opposed to swapping a mutated prompt for a previous task prompt.
Holistically, our work does not replace previous methods but instead utilizes them as operators in a pipeline to automatically construct novel data with minimal human involvement and trivial amounts of seed data. This is due to features of massive parallel scaling, operator neutrality and compound mutations.
@@ 3 Genetic-Instruct
We introduce Genetic-Instruct, an algorithm inspired by the population-based evolutionary approach of genetic algorithms. This algorithm employs two primary evolutionary operations of mutation and crossover to evolve and generate new generations from an initial population. The initial population, termed Generation 0, comprises a limited set of high-quality seed instructions. These seed instructions undergo a series of evolutionary operations, mainly mutation, crossover and selection, to transform them into new instructions. In this study, the operations of mutation, crossover, and selection are executed by leveraging large language models (LLMs) and enhancing their output with in-context learning (Brown et al., 2020).
We designate the LLM responsible for instruction generation as the Instructor-LLM. Upon generating a new instruction, another LLM, referred to as the Coder-LLM, is tasked with producing the code corresponding to this new instruction. The Instructor-LLM and Coder-LLM can either be distinct models or the same model but with different prompts. The newly generated instruction and its associated code constitute a new coding instruction, which can be utilized for training. However, there may be instances where the generated code does not fully address the provided question, or the question itself may be poorly formulated. To assess the quality of the new coding instruction, we employ another LLM, termed the Judge-LLM, to evaluate the correctness of the instruction and its code. If a sample passes this quality assessment, it is added to the pool of instructions and may be selected as the seed instruction for the next batch of synthesized samples.
The entire process is iterated multiple times to synthesize samples until the desired population size is achieved. Subsequently, a deduplication process is applied to minimize duplicate instructions in the current dataset. This resulting population is then labeled as a generation, and the entire pipeline can be repeated by considering this generation as the initial population for the next generation. The complete pipeline is illustrated in Figure 1, and the procedure for generating one generation is detailed in Algorithm 9.
3.1 Selection Process
Initially, from the instruction set of the initial population ($S$) (also called the seed dataset), we randomly select a batch of instructions ($S^{\prime}$) with replacement. This batch is then subjected to either mutation or crossover steps as detailed below, with the generations denoted by $S^{\prime}_{gen}$. In subsequent iterations, we augment the initial population with the generations from the previous iteration, such that $S:=S\cup S^{\prime}_{gen}$. This process continues until a sufficient number of samples have been synthesized for the current generation.
@@ 3.2 Mutation Operation
The mutation operation is inspired by an adaptation of the Evol-Instruct algorithm, as devised by Xu et al. (2024), and further extended by WizardCoder (Luo et al., 2024) to facilitate instruction generation for code models. Evol-Instruct evolves an instruction into another using an LLM based on predefined tasks. For a sample selected for mutation, we randomly choose one of the five tasks defined and apply the mutation to generate a new instruction. We employ the same five tasks introduced by Luo et al. (2024), with minor prompt modifications to suit our Instructor-LLM. Details on the mutation prompts are provided in Figure 2.
@@ 3.3 Crossover Operation
The crossover operation in Genetic-Instruct is influenced by concepts introduced in Self-Instruct (Wang et al., 2023) and Unnatural Instructions (Honovich et al., 2023). It combines multiple genes (instructions) and employs the Instructor-LLM to generate new populations from the provided few-shot example instructions. To enhance the efficiency of the crossover operation, we provide multiple seed instructions and request the model to generate multiple diverse new instructions based on the provided examples in a single Instructor-LLM call. Simple post-processing, such as rejecting all samples that don’t pass the Abstract Syntax Tree checks, is then applied to filter out incorrect instructions. The prompt for the crossover operation is depicted in Figure 3.
@@ 3.4 Code Generation
After the Instructor-LLM generates a batch of new instructions, they are passed to the Coder-LLM to generate the corresponding code solutions. The Coder-LLM should be proficient in coding tasks to ensure the generation of high-quality solutions. However, some generated code may not be parseable or compilable. Therefore, we filter out solutions whose code segments cannot be parsed by the corresponding language’s parser/compiler. While determining the correctness of code by execution is the ideal case, it is challenging due to various factors, such as language constraints, missing dependencies, or having to integrate the current solution into a much larger codebase that may not be available in its entirety. The prompt used in this step is illustrated in Figure 4.
@@ 3.5 Fitness Function
In genetic algorithms, once candidate solutions are produced, they are scored using a fitness function in order to discard candidates that score lower compared to others. Similarly, we employ a Judge-LLM to assign a binary score indicating whether a candidate code solution meets the requirements of the corresponding instruction. The Judge-LLM is provided with an instruction and its code solution to determine the correctness of the instruction and its corresponding solution. To enhance the Judge-LLM’s performance, we employ techniques such as in-context learning with few-shot examples and Chain-of-Thought (Wei et al., 2022) prompting to facilitate reasoning before making a binary decision. The prompt for the Judge-LLM is depicted in Figure 5.
@@ 3.6 Scaling Up the Process
An advantage of genetic algorithms is their inherent capacity for parallelization. When utilizing computationally intensive LLMs for sample generation, it is crucial to exploit this parallel architecture. We efficiently compute multiple colonies of populations in parallel processes or compute nodes and synchronize them periodically. These colonies are evolved and populated independently, starting from the same seed population. Upon reaching the desired size, the colonies are merged into a single population, where deduplication is performed to obtain a unified dataset, termed one generation. We use only the instruction itself as the deduplication criterion to eliminate redundancies across different colonies. Deduplication is performed using the MinHash + LSH algorithm, implemented in text-dedup library. This generation serves as the shared seed data for the next generation, a process we refer to as inter-node parallelization.
Within each colony, using large batch sizes and selecting a single task for each algorithm iteration leads to redundancy in the generated instructions, and increases duplication. To address this, we employ sub-processes to compute multiple sub-colonies within each colony, each possessing a small batch size but capable of selecting diverse tasks. This approach maintains similar computational and memory requirements as a single large batch of single-task samples while enhancing the diversity of generated solutions, thereby reducing duplication. This process is termed intra-node parallelization.
@@ 4 Experiments
In all experiments, the models are assessed using four benchmark datasets: HumanEval (Chen et al., 2021), MBPP (Odena et al., 2021), HumanEval+, and MBPP+ (Liu et al., 2023). The MBPP+ and HumanEval+ datasets from the EvalPlus benchmark are extensions of the MBPP and HumanEval test sets, respectively. These extensions are enriched with additional test cases to ensure the correctness and accuracy of the generated code. All coding evaluations are conducted using greedy decoding. In all reported results, an asterisk (*) indicates that we computed the results ourselves, as the original works did not report these numbers.
We use the Tiger-Leetcode collection [^1] TigerResearch/tigerbot-kaggle-leetcodesolutions-en-2k as the initial population in most of the experiments. This collection serves as the seed dataset for the generation process and contains interview-style coding questions and solutions.
In all the experiments, we employed the same generation model as Instructor-LLM, Coder-LLM, and Judge-LLM. Since we are evaluating our approach solely on Python coding benchmarks, we explicitly limit our generated solutions to Python by requesting the models to generate only Python code. After the code is generated by Coder-LLM, we verify its syntactical correctness using Python’s ast package, regardless of its executability.

### v2

@@ Abstract
Large Language Models (LLMs) require high quality instruction data for effective alignment, particularly in code generation tasks where expert curated datasets are expensive to produce. We present Genetic-Instruct, a scalable algorithm for synthesizing large-scale, high quality coding instructions using evolutionary principles. Starting from a small set of seed instructions, Genetic-Instruct generates diverse and challenging instruction-code pairs by leveraging an Instructor-LLM for generation, a Coder-LLM for code synthesis, and a Judge-LLM for automatic quality evaluation. Our proposed approach is highly parallelizable and effective even with a small seed data and weaker generator models. We generated more than 7.5 million coding instructions with the proposed approach. Then we evaluated it by fine-tuning LLMs with the synthetic samples and demonstrated a significant improvement in their code generation capability compared to the other synthetic generation approaches and publicly available datasets. Our results highlight the efficiency, scalability, and generalizability of the Genetic-Instruct framework.
Genetic Instruct: Scaling up Synthetic Generation of Coding Instructions for Large Language Models
Somshubra Majumdar*, Vahid Noroozi thanks: Equal contribution, Mehrzad Samadi, Sean Narenthiran, Aleksander Ficek, Wasi Uddin Ahmad, Jocelyn Huang, Jagadeesh Balam, Boris Ginsburg NVIDIA Santa Clara, CA 15213, USA {smajumdar,vnoroozi,msamadi,snarenthiran,aficek, wasiuddina,jocelynh,jbalam,bginsburg}@nvidia.com
@@ 1 Introduction
Large Language Models (LLMs) have made significant progress in programming tasks and are increasingly used as code assistants (Liang et al., 2024). To fully exploit their potential, they require alignment (Ouyang et al., 2022), which depends on paired instruction-solution examples to shape the behavior of the model. However, creating diverse and complex instructions, especially in coding domains, can be expensive due to the need for expert input. A promising alternative is generating synthetic instructions using another LLM. Prior research shows that synthetic instructions are effective for both coding (Luo et al., 2024; Wu et al., 2024; Wei et al., 2024b; Yu et al., 2024) and general tasks (Wang et al., 2023; Honovich et al., 2023; Xu et al., 2024).
@@
Figure 1: The overall process of Genetic-Instruct across multiple parallel colonies per generation. Each colony begins with a small seed population, from which an Instructor-LLM applies crossover and mutation to create new instructions. A Coder-LLM then generates corresponding code solutions, which are evaluated by a Judge-LLM for correctness and quality. Once the target population size is reached, samples are decontaminated to form the final population.
In this paper, we introduce Genetic-Instruct, a scalable algorithm for generating synthetic coding instructions, illustrated in Figure 1. Inspired by evolutionary algorithms, Genetic-Instruct starts with a small set of seed instructions and uses LLMs to generate new instruction-code pairs through two operations of crossover and mutation.
The crossover operation follows a self-instruct approach Wang et al. (2023), where an LLM creates new instructions from few-shot examples, expanding topic coverage beyond the original seeds. The crossover operator is mainly employed to enhance diversity by expanding the overall coverage of the instructions to wider domains and topics beyond the original seed instructions.
In the mutation operation, an LLM evolves a given instruction into another instruction based on some predefined rules (Luo et al., 2024). This operation can help the generation process to increase the diversity of the instructions locally. An instruction generated by one operation is added to the pool of the seeds and it may be be used by the the operation or other in the next step. This collaborative and coupled interaction between the crossover and mutation is the main key foundation of our proposed approach. It boosts instruction diversity which is an essential factor in the success of synthetic instruction generation.
Subsequently, another LLM generates answers, including code solutions, for the instructions. We introduce a fitness function that uses an LLM to evaluate the correctness and quality of each instruction-solution pair. Samples that pass these checks are added to the population pool, and the evolutionary process continues until the target population size is reached. Starting from a small set of seed instructions, the pool grows with newly generated synthetic instructions.
Additionally, the entire pipeline is designed for efficient parallel execution with multiple colonies of populations by running multiple instances of this process in parallel. Furthermore, this process can get repeated multiple times to generate more generations by using the instructions generated from previous round as the seed for the next generation.
Using our Genetic-Instruct algorithm, we generated a large dataset of synthetic coding instructions (more than 7.5M samples), starting from 512 seed questions. We trained LLMs on this data via supervised fine-tuning (SFT) and evaluated them on code generation benchmarks. Our work supports open-source development, avoiding any closed-source data or models.
Models trained on our synthetic dataset achieved strong results across coding benchmarks, outperforming other instruction generation methods and also some of the existing public SFT datasets. Our experiments also show that Genetic-Instruct can produce high-quality data without requiring very strong LLMs or large seed sets.
@@ 2 Previous Works
Synthetic data generation has become a practical alternative to the costly and time-consuming collection of human-curated data for LLM training. A notable method is Self-Instruct (Wang et al., 2023), which uses a pre-trained LLM to generate instruction-output pairs from a small seed set, then fine-tunes the base model. However, Self-Instruct focuses on general tasks, not coding. Moreover, while it can enhance the coverage of topics, the synthesized samples are often simple and not challenging enough to require additional steps to arrive at the solution.
To overcome this, Evol-Instruct Xu et al. (2024) introduces instruction mutation to create more complex and diverse tasks through meta-instructions that increase reasoning depth, impose constraints, or promote conceptual evolution. This idea was adapted to coding by WizardCoder Luo et al. (2024), leading to improved coding performance in models trained on such evolved instructions.
While Self-Instruct and Evol-Instruct generate instructions without using any code as seeds, another line of work Yu et al. (2024); Wu et al. (2024); Wei et al. (2024b) generates instructions from existing code snippets. These approaches leverage large code corpora to synthesize diverse prompts. For example, INVERSE-CODER Wu et al. (2024) generates instructions directly matched to given code, whereas OSS-Instruct Wei et al. (2024b) and WaveCoder Yu et al. (2024) use LLMs to create new, code-inspired instructions. However, these methods rely on large high quality and processed code samples, which may pose challenges for less common programming languages.
Input :  $N$: Number of colonies
$P_{max}$: Maximum population size per colony
$G_{N}$: Total number of generations
$B_{m}$ and $B_{c}$: Number of individuals needed for mutation and cross-over respectively
$P_{seed}$: Initial set of seed instructions
$M_{p}$: Probability of selecting mutation as operator
$P_{op}$: Probability distribution over the operations {Mutation: $M_{p}$, Cross-over: $1-M_{p}$}
Output : $FinalInstructions$: Generated Synthetic Instructions for Coding Problems
for $g\leftarrow 1$ to $G_{N}$ do
       Run $N$ colonies in parallel;
       foreach colony do
             Initialize $P_{pool}\leftarrow P_{seed}$;
             while len($P_{pool}$) $<$ $P_{max}$ do
                   $OP\leftarrow$ Choose an operation from $P_{op}$;
                   $Candidates\leftarrow$ Select a subset of $B_{m}$ or $B_{c}$ individuals from $P_{seed}$ randomly based on the selected operation;
                   $NewQuestions\leftarrow InstructorLLM(Candidates,OP)$;
                   $FilteredQuestions\leftarrow FilterQuestions(NewQuestions)$;
                   $GeneratedInstructions\leftarrow CoderLLM(FilteredQuestions)$;
                   $ValidatedInstructions\leftarrow ValidateCode(GeneratedInstructions)$;
                   $NewInstructions\leftarrow JudgeLLM(ValidatedInstructions)$;
                   $P_{pool}\leftarrow P_{pool}\cup NewInstructions$;
             end while
       end foreach
      $G_{g}\leftarrow$ Aggregate all $P_{pool}$ from $N$ colonies;
end for
$AggInstructions\leftarrow$ Aggregate all $G_{g}$, for $g\in[1,G_{n}]$;
$FinalInstructions\leftarrow Decontaminate(AggInstructions)$;
Algorithm 1 Pseudo-code for the Genetic-Instruct Algorithm
@@ 3 Genetic-Instruct
We introduce Genetic-Instruct, an algorithm inspired by the population-based genetic algorithms Golberg (1989). This algorithm employs the two primary evolutionary operations of mutation and crossover to evolve and generate new generations from an initial population. The initial population, termed Generation 0, comprises a limited set of high-quality seed instructions. These seed instructions undergo a series of evolutionary operations, mainly mutation, crossover and selection, to transform them into new instructions. All the operations are executed by leveraging LLMs and enhancing their output with in-context learning.
The whole process of Genetic-instruct is as follows. At each step, from the instruction set of the initial population (seed population), we randomly select a batch of instructions with replacement. The LLM responsible for instruction generation (called Instructor-LLM) is employed to synthetize the new instructions based on a selected operation. Upon generating a new instruction, another LLM, referred to as the Coder-LLM, is tasked with producing the code corresponding to this new instruction. The newly generated instruction and its associated code constitute a new coding instruction, which can be utilized for training. However, there may be instances where the generated code does not fully address the provided question, or the question itself may be poorly formulated. To assess the quality of the new coding instruction, we employ another LLM, termed the Judge-LLM, to evaluate the correctness of the instruction and its code. If a sample passes this quality assessment, it is added to the pool of instructions and may be selected as the seed instruction for the next batch of synthesized samples. The entire process is iterated multiple times to synthesize samples until the desired population size is achieved. This resulting population is then labeled as a generation, and the entire pipeline can be repeated by considering this generation as the initial population for the next generation.
Subsequently, a decontamination process is applied to minimize risk of contaminated instructions in the training data. The complete pipeline is illustrated in Figure 1 for one generation, and the procedure for the whole algorithm is detailed in Algorithm 1. In the following section, each step is explained in detail.
@@ 3.1 Mutation Operation
The mutation operation is inspired by an adaptation of the Evol-Instruct algorithm, as devised by Xu et al. (2024), and further extended by WizardCoder (Luo et al., 2024) to facilitate instruction generation for code models. Evol-Instruct evolves an instruction into another using an LLM based on predefined tasks. For a sample selected for mutation, we randomly choose one of the five tasks defined and apply the mutation to generate a new instruction. We employ the same five tasks introduced by Luo et al. (2024), with minor prompt modifications to suit our Instructor-LLM. Details on the mutation prompts are provided in Appendix 3.
@@ 3.2 Crossover Operation
The crossover operation in Genetic-Instruct is influenced by the concepts introduced in Self-Instruct (Wang et al., 2023) and Unnatural Instructions (Honovich et al., 2023). It inspires from multiple instructions and employs the Instructor-LLM to generate new populations from the provided few-shot example instructions. To enhance the efficiency of the crossover operation, we provide multiple seed instructions and request the model to generate multiple diverse new instructions based on the provided examples in a single Instructor-LLM call. The prompt for the crossover operation is depicted in Appendix 4.
@@ 3.3 Code Generation
After the Instructor-LLM generates a batch of new instructions, they are passed to the Coder-LLM to generate the corresponding code solutions. The Coder-LLM should be proficient in coding tasks to ensure the generation of high-quality solutions. However, some generated code may not be parseable or compilable. Therefore, we filter out solutions whose code segments cannot be parsed by the corresponding language’s parser/compiler. While determining the correctness of code by execution is the ideal case, it is challenging due to various factors, such as language constraints, missing dependencies, or having to integrate the current solution into a much larger codebase that may not be available in its entirety. The prompt used in this step is illustrated in Appendix 5.
@@ 3.4 Fitness Function
Simple post-processing, such as rejecting all samples that don’t pass the Abstract Syntax Tree checks, is applied to filter out incorrect instructions. Then, they are scored using a fitness function in order to discard candidates that have low quality. We employ a Judge-LLM to assign a binary score indicating whether a candidate code solution meets the minimum requirements. The Judge-LLM is provided with an instruction and its code solution to determine the correctness of the instruction and its corresponding solution. To enhance the performance, we employ techniques such as in-context learning with few-shot examples and Chain-of-Thought (Wei et al., 2022) prompting to making a better decision. The prompt for the Judge-LLM is depicted in Appendix 6.
@@ 3.5 Scaling Up the Process
An advantage of genetic algorithms is their inherent capacity for parallelization. When utilizing computationally intensive LLMs for sample generation, it is crucial to leverage this parallel structure. We execute multiple colonies of populations in parallel processes and synchronize them periodically. These colonies are evolved and populated independently, starting from the same seed population. Upon reaching the desired size, the colonies are merged into a single population and called a generation. Additionally, to improve the diversity, we make sure that seed examples selected to be used in a batch are all different.
3.6 LLM Decontamination
To prevent any evaluation benchmark questions from leaking into our training samples, we adopted the decontamination methodology proposed by Yang et al. (2023), which involves two primary stages. First, for each synthesized question, we performed an embedding-based similarity search using a Sentence Transformer Reimers and Gurevych (2020) model to identify the most similar test example from all benchmark datasets. Second, we constructed question pairs by matching each synthesized question with its most similar test example. An LLM, specifically Meta-Llama-3-70B-Instruct, was then employed to evaluate whether any of these pairs constituted a paraphrase (details on the prompt are provided in Appendix  7). To control for potential positional bias in the LLM’s paraphrase detection, we generated two pairs for each match: one where the synthesized question appeared first and another where the test set question was presented first (Toshniwal et al., 2024). If any of these pairs were determined to be similar by the LLM, the synthesized question was removed.
@@ 4 Experiments
We fine-tune the base LLM models using supervised fine-tuning (SFT) to evaluate the effectiveness of a given instruction set. In all experiments, the models are evaluated on four benchmark datasets: HumanEval (HE) (Chen et al., 2021), MBPP (Odena et al., 2021), HumanEval+ (HE+), and MBPP+ (Liu et al., 2023). The MBPP+ and HumanEval+ datasets, part of the EvalPlus benchmark, are extensions of the original MBPP and HumanEval test sets, respectively. These extensions include additional test cases designed to ensure the correctness and accuracy of the generated code. The prompts used for the evaluation benchmarks are provided in Appendix 9. All code evaluations are conducted using greedy decoding. Prior to SFT training, all training datasets undergo a decontamination process.
We use 512 samples from the Tiger-Leetcode collection (TigerResearch, 2023) as the initial population in most experiments. This collection serves as the seed dataset for the first generation and consists of interview-style coding questions. Throughout all experiments, we employ the same generation models as Instructor-LLM, Coder-LLM, and Judge-LLM. Since our evaluation focuses exclusively on Python coding benchmarks, we constrain the generated solutions to Python by instructing the models to produce only questions that can be answered with Python code. After code is generated by Coder-LLM, we verify its syntactic correctness using Python’s ast package, regardless of its executability, to ensure the structural validity of the generated code.